What is the maximum number of electrons that can be emitted if a potassium metal surface absorbs 3.50✖️10^-3 J of radiation at a wavelength of 250.0 nm? The work function (binding energy) of potassium metal is 3.845✖️10^-19 J. What is the velocity of the elctron emitted?
Answer #1

#(a):

Given wavelength of absorbed radiation, = 250.0 nm * (10-9 m / 1 nm) = 2.50*10-7 m

=> Frequency of the absorbed light, = C / = 3*108 m.s-1 / 2.50*10-7 m = 1.2*1015 Hz or s-1  

Suppose there are "n" photons in the radiation of 3.50*10-3 J.

=> energy of a single absorbed photon, E = h* = 6.626*10-34 J.s * 1.2*1015 s-1 = 7.95*10-19 J

"n" number of photons will also produce "n" number of electrons

=> maximum number of electrons formed by 3.50*10-3 J energy = Total energy / energy of single photon

= 3.50*10-3 J / 7.95*10-19 J = 4.40*1015 electrons (Answer)

#(b): energy of a single absorbed photon, E = 7.95*10-19 J

work function, = 3.845*10-19 J

=> Amount of absorbed energy that converts to kinetic energy of free electron, Ee = E - =  7.95*10-19 J -  3.845*10-19 J

= 4.105*10-19 J

Also Ee = (1/2)*me * V2

where me = mass of electron = 9.11*10-31 Kg

=> V = square root(2Ee / me)

=> V = square root(2*4.105*10-19 J / 9.11*10-31 Kg )  

=> V = 9.49*105 m/s (Answer)